# Teaching Spell Checkers to Teach: Pedagogical Program Synthesis for Interactive Learning

### v1

@@ 9. Conclusion
This work addressed a fundamental gap in literacy instruction: while speech-language pathologists employ inquiry-based approaches to teach spelling through structured word investigation, everyday writing tools remain limited to detection and autocorrection. We introduced SPIRE (Spelling Inquiry Engine), which brings structured word inquiry into the act of composition through Pedagogical Program Synthesis, a novel approach that operationalizes the inherently dynamic pedagogy of spelling instruction. Central to our contribution is a domain-specific language that represents SLP instructional moves as composable primitives, enabling the system to synthesize tailored programs in real-time from learner errors and render them as interactive interfaces. Each synthesized program guides learners through brief cycles of hypothesis testing on word meanings, structures, morphological families, origins, and grapheme-phoneme correspondences, transforming spelling errors into opportunities for metalinguistic reasoning alongside correction. Evaluation with five speech-language pathology experts demonstrated strong pedagogical validity, with near-ceiling ratings on reasoning appropriateness and high interrater agreement. A preliminary usability assessment with seven children revealed positive engagement and identification of morphological investigation as a novel and meaningful learning experience absent from current classroom instruction. SPIRE demonstrates that program synthesis can capture expert pedagogical reasoning that purely rule-based systems cannot, offering a vessel for structured literacy experiences at scale. This work suggests broader applicability of Pedagogical Program Synthesis to instructional domains requiring flexible, context-sensitive reasoning about learner understanding, extending beyond spelling to other areas where expert practice must adapt dynamically to individual learner needs.

### v2

@@ 9. Conclusion
This work addressed a fundamental gap in literacy instruction: while speech-language pathologists employ inquiry-based approaches to teach spelling through structured word investigation, everyday writing tools remain limited to detection and autocorrection. We introduced SPIRE (Spelling Inquiry Engine), which brings structured word inquiry into the act of composition through Pedagogical Program Synthesis, a novel approach that operationalizes the inherently dynamic pedagogy of spelling instruction. Central to our contribution is a domain-specific language that represents SLP instructional moves as composable primitives, enabling the system to synthesize tailored programs in real-time from learner errors and render them as interactive interfaces. Each synthesized program guides learners through brief cycles of hypothesis testing on word meanings, structures, morphological families, origins, and grapheme-phoneme correspondences, transforming spelling errors into opportunities for metalinguistic reasoning alongside correction. Evaluation with ten speech-language pathology experts demonstrated strong pedagogical validity, with near-ceiling ratings on reasoning appropriateness and high interrater agreement. A preliminary usability assessment with seven children revealed positive engagement and identification of morphological investigation as a novel and meaningful learning experience absent from current classroom instruction. SPIRE demonstrates that program synthesis can capture expert pedagogical reasoning that purely rule-based systems cannot, offering a vessel for structured literacy experiences at scale. This work suggests broader applicability of Pedagogical Program Synthesis to instructional domains requiring flexible, context-sensitive reasoning about learner understanding, extending beyond spelling to other areas where expert practice must adapt dynamically to individual learner needs.
@@ 10. GenAI Usage Disclosure
Acknowledgements.
We are grateful to the reviewers and our study participants for their time and helpful feedback. This work is supported through the AI Research Institutes program by the National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award #2229873\#2229873 - National AI Institute for Exceptional Education.